OpenStudy (anonymous):

I need help with this question. 3x=4+4y.

OpenStudy (anonymous):

What about it?

OpenStudy (anonymous):

How do you solve it with using elimination?

OpenStudy (anonymous):

You need two equations.

OpenStudy (anonymous):

5x+4y=12 and 3x=4+4y

OpenStudy (anonymous):

Arrange the equations to the same format. x + y = #

OpenStudy (anonymous):

5x+ 4y=12 3x - 4y=4 ------------ 5x+3x = 8x 4y-4y=0 12+4=16 8x=16 x=2 Then plug the 2 in for x in one of the equations and solve for y.